you are buying scarves and gloves as gifts. a scarf costs $2 less than each pair of gloves. gloves cost $7. you buy 12 items for $62. how many gloves did you buy ?
Found 2 solutions by nerdybill, ewatrrr:
Answer by nerdybill(7384)

you are buying scarves and gloves as gifts. a scarf costs $2 less than each pair of gloves. gloves cost $7. you buy 12 items for $62. how many gloves did you buy ?
.
Since
g = pairs of gloves purchased
then
12-g = scarfs purchased
.
7g + 5(12-g) = 62
7g + 60-5g = 62
2g + 60 = 62
2g = 2
g = 1 (gloves purchased)

Answer by ewatrrr(24785)


Hi,
Let x and (12-c) represent the number of gloves and scarves respectively
Note: a scarf cost $5 ($2 less than gloves cost)
Question states
$7x + $5(12-x) = $62
2x = $2.00
x = 1 , the number of gloves. Number of scarves is 11. (12-1)
CHECKING our Answer
$7*1 + $5 * 11 = $7 + $55 = $62